OpenStudy (anonymous):

A generous man set aside a certain sum of money for equal distribution weekly to the needy of his acquaintance. One day he remarked,"If there are five fewer applicants next week, you will each receive two dollars more." Unfortunately, Instead of there being fewer there were actually four more persons applying for the gift. "This means," he pointed out, "that you will each receive one dollar less." How much did each person receive at that last distribution?

OpenStudy (anonymous):

okay, first let's have \(n\) be the number of people and \(t\) be the total money. So what equations can we make?

OpenStudy (anonymous):

I mean \(n\) will be the number of people there were before the four extra came.

OpenStudy (anonymous):

So let's acknowledge that the average would have been \(\frac t n \) had it stayed the same. ""If there are five fewer applicants next week, you will each receive two dollars more"\[ \frac{t}{n-5} = \frac t n +2 \]"actually four more persons applying for the gift. "This means," he pointed out, "that you will each receive one dollar less.""\[ \frac{t}{n+4} = \frac t n -1 \]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

WE can combine t hose two equations: \[ \frac{t}{n-5} - 2 = \frac t {n+4} +1 \]

OpenStudy (anonymous):

We can do lots of things. You should use your algebra skill to win.

OpenStudy (anonymous):

okay ^^

OpenStudy (anonymous):

is there another way to do this without using algebra?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."